# On the failure of ReLU activation for physics-informed machine learning

### v1

@@ 1 Introduction
The physics here are beside the point—it suffices to say that $x$ is the spatial coordinate, $u(x)$ is the solution field, $b(x)$ is a source term, and $\kappa(x)$ is a coefficient field reminiscent of a material property. We call this the “strong form" of the governing equation. As is well-known in the engineering mechanics community [2], obtaining a solution to Eq. (1) can be equivalently thought of as the minimization of a particular “energy" functional:
@@
The calculus of variations can be used to show that the minimizer of the variational energy $\Pi$ satisfies an Euler-Lagrange equation given by Eq. (1) [11]. We call Eq. (2) the “variational form" of the problem. Regardless of the perspective we take on the solution, the general lack of analytical solutions to Eqs. (1) and (2) necessitates that we resort to approximations. Most traditional numerical methods represent the unknown solution field in a basis $\{f_{i}(x)\}_{i=1}^{N}$, where $N$ is the number of basis functions and each $f_{i}$ is chosen by the analyst. The representation of the solution is thus $u(x)=\sum_{i=1}^{N}\theta_{i}f_{i}(x)$, where $\boldsymbol{\theta}$ are the unknown degrees of freedom which specify the approximation. By way of example, spectral methods [10] and finite element methods [14] are two classical strategies which differ only in their choice of basis.
@@
where $\mathbf{x}_{k}$ is the input to the $k$-th layer, $\mathbf{A}_{k}$ are the weights, and $\mathbf{b}_{k}$ are the biases. The output $\mathbf{y}_{k}$ becomes the input to the next layer $\mathbf{y}_{k}\rightarrow\mathbf{x}_{k+1}$. The function $\sigma(\cdot)$ is a nonlinear activation function that is applied element-wise. To solve differential equations with the standard PINNs approach, an optimization objective is formed by taking the strong form of Eq. (1), plugging in the neural network discretization, squaring the discretized differential equation pointwise, and integrating it over the domain. The parameters are then chosen such that this objective is minimized. Whereas boundary conditions were originally treated with additional penalties in the objective, an elegant alternative for the case of Dirichlet boundaries is given by [23], where the boundary conditions are enforced automatically by multiplying the neural network with a “distance function" $D(x)$. For our one-dimensional problem with homogeneous Dirichlet boundaries, the solution is thus represented as
@@
where $\sin(\pi x)$ is the distance function. This same discretization can be used in the “Deep Ritz" formulation of the problem [8], which obtains a numerical solution through minimizing the variational energy of Eq. (2) in terms of the neural network parameters.
Whether physics-informed or not, the choice of activation function plays a fundamental role in the training and performance of a neural network. Some common choices of activation functions are hyperbolic tangent ($\tanh(\bullet)$), sigmoid ($1/(1+\exp(-\bullet))$), softplus ($\ln(1+\exp(\bullet))$), and the rectified linear unit (ReLU) ($\max(0,\bullet)$). Other options include the swish activation and the Gaussian error linear unit (GELU) [12]. The problem of choosing activation functions has been studied extensively in the physics-informed machine learning literature. For example, in [27], using non-standard “hybrid" activation functions was shown to reduce errors in PINN solutions. Similarly, [15] showed that allowing the activation functions themselves to be partially learnable accelerated convergence in the physics-informed training. Another study found that learning the activation function allowed the network to better adapt to characteristics of the solution field [25]. For high-frequency solution fields, modifications inspired by the SIREN activation [21] were shown to reduce the number of parameters and thus expedite training [9]. In a comparative study on the Helmholtz equation, the swish activation function outperformed other activation functions by obtaining the smallest error with a reference solution [1]. In another work, the one-dimensional wave equation was used to compare the convergence and errors obtained from different activation functions, including hyperbolic tangent, swish, and sinusoidal activation functions, finding that no single activation performed best in all numerical experiments [17]. The authors remark that their choice to use the strong form of the governing equation places restrictions on the set of available activation functions. Namely, the governing equation requires second derivatives of the solution field, and the second derivative of ReLU is zero everywhere except at the origin. As such, they do not include ReLU in their study. Finally, [7] compared six different activation functions including hyperbolic tangent, ReLU, swish, and modifications thereof on incompressible flow problems, finding that ReLU consistently performed the worst.
@@
We use ADAM optimization with a learning rate of $1\times 10^{-3}$ to minimize the variational energy. Integration is done on a uniform grid of $250$ points and we run the optimization for $5000$ epochs. The exact solution is obtained by inspection as $\frac{100}{9\pi^{2}}\sin(3\pi x)$. As an additional point of reference, we train a second single hidden-layer network of the same width with hyperbolic tangent activation functions. See Figure 1 for the results. The converged energy of the ReLU network is notably larger than that of the network using hyperbolic tangent activation functions. The eyeball norm suffices to show that the ReLU network fails to reproduce the simple solution field. As the figure shows, this failure occurs in spite of the fact that the network can represent the exact solution when explicitly trained to do so. Figure 2 shows similar results for the higher-frequency source term $b(x)=100\sin(4\pi x)$. Evidently, these failures are not issues of expressiveness. We remark that this jagged shape of the solution field is similar to results obtained in [7], where ReLU provides very poor approximations of solution fields from incompressible flow problems. If the ReLU network is expressive enough to represent the true solution, and it possesses the same continuity properties as finite element discretizations, why does it perform so poorly with the Deep Ritz method? This is the question we address in the subsequent section.
@@
Figure 2: Training convergence of the variational energy objective with a higher-frequency source term for the two networks (left). The Deep Ritz solution with ReLU activation again provides an unphysical and inaccurate representation of the solution (center). We verify that the network is sufficiently expressive to represent the solution field (right).
@@ 2 Diagnosing the failure of ReLU in the Deep Ritz problem
where the discretized solution is given by Eq. (4). This gradient is used in the training process to iteratively update the parameters, but is often “out of sight" for the analyst, as it is computed by automatically differentiating the forward evaluation of the energy. We now compute each of the terms involved in Eq. (5) analytically. Recall that the solution field is given by the single hidden-layer network as
@@
Notice that second derivatives of ReLU appear in Eq. (8). This is a consequence of the nonlinearity of the discretization—while the $C^{0}$ continuity of the solution field is appropriate for a traditional Ritz method, training the network with the Deep Ritz method requires second derivatives of the activation. This is because trainable parameters lie within the ReLU activation, unlike a so called “linear" discretization, where the parameters only multiply spatial basis functions. Mathematically speaking, the delta function arising from the second derivative of ReLU is acceptable given that it appears under the integral in the variational energy. However, it is not immediately clear how PyTorch’s automatic differentiation handles gradients of discontinuous functions. We will investigate this numerically. To do this, we compute derivatives of the function $\sin(\pi x)\max(0,x-1/2)$ both analytically and with automatic differentiation. This function is reminiscent of one of the $N$ learned “basis" functions that are scaled and superimposed in the ReLU network of Eq. (4). To distinguish between the two cases of analytic and automatic differentiation, we call the analytically differentiated function $u(x)$ and the automatically differentiated function $w(x)$. Note that we can use Eqs. (6)-(8) with $N=1$, $\theta^{1}=1$, $\theta^{2}=-1/2$, and $\theta^{3}=1$ to obtain the relevant derivatives. To numerically approximate the delta function on a fixed integration grid, we require that
@@
where the notation $\mathbf{1}_{i}$ indicates a vector of zeros with a single $1$ in the $i$-th index. We take the finite difference step size to be $h=1\times 10^{-3}$, the learning rate to be $5\times 10^{-3}$, and run ADAM optimization for $7500$ epochs. The first single hidden-layer ReLU network is trained with gradients from automatic differentiation (AD) and the second is trained with manually supplied gradients computed using finite differences (FD). See Figure 6 for the results of this comparison. The network with finite differenced gradients obtains a lower converged energy value, and a smoother, more accurate solution field.
@@ 3 Conclusion
The purpose of this short work has been to explain failures of ReLU activation functions for physics-informed machine learning problems. Though the objective in the Deep Ritz test problem we considered required only first derivatives of the solution field, we showed that training the network requires second derivatives of the activation function. The second derivatives of the $C^{0}$ ReLU activation give rise to delta functions which automatic differentiation does not account for, thus leading to inaccurate gradient computations. We obtained accurate gradient computations by finite differencing the loss function, and showed that this remedied the failures of the ReLU network on the test problem. Though the corrected gradients improved performance, we then showed that the loss landscape of our test problem with the ReLU network is extremely noisy, and that this noise promptly vanishes when higher-order continuity is introduced. This study suggests that ReLU activation is a poor choice for physics-informed machine learning, which corroborates the conclusions of past comparisons of activation functions in the PINNs literature.

### v2

@@ 1 Introduction
The physics here are beside the point—it suffices to say that $x$ is the spatial coordinate, $u(x)$ is the solution field, $b(x)$ is a source term, and $\kappa(x)$ is a coefficient field reminiscent of a material property. We call this the “strong form” of the governing equation. As is well-known in the engineering mechanics community [2], obtaining a solution to Eq. (1) can be equivalently thought of as the minimization of a particular “energy” functional:
@@
The calculus of variations can be used to show that the minimizer of the variational energy $\Pi$ satisfies an Euler-Lagrange equation given by Eq. (1) [11]. We call Eq. (2) the “variational form” of the problem. Regardless of the perspective we take on the solution, the general lack of analytical solutions to Eqs. (1) and (2) necessitates that we resort to approximations. Most traditional numerical methods represent the unknown solution field in a basis $\{f_{i}(x)\}_{i=1}^{N}$, where $N$ is the number of basis functions and each $f_{i}$ is chosen by the analyst. The representation of the solution is thus $u(x)=\sum_{i=1}^{N}\theta_{i}f_{i}(x)$, where $\boldsymbol{\theta}$ are the unknown degrees of freedom which specify the approximation. By way of example, spectral methods [10] and finite element methods [14] are two classical strategies which differ only in their choice of basis.
@@
where $\mathbf{x}_{k}$ is the input to the $k$-th layer, $\mathbf{A}_{k}$ are the weights, and $\mathbf{b}_{k}$ are the biases. The output $\mathbf{y}_{k}$ becomes the input to the next layer $\mathbf{y}_{k}\rightarrow\mathbf{x}_{k+1}$. The function $\sigma(\cdot)$ is a nonlinear activation function that is applied element-wise. To solve differential equations with the standard PINNs approach, an optimization objective is formed by taking the strong form of Eq. (1), plugging in the neural network discretization, squaring the discretized differential equation pointwise, and integrating it over the domain. The parameters are then chosen such that this objective is minimized. Whereas boundary conditions were originally treated with additional penalties in the objective, an elegant alternative for the case of Dirichlet boundaries is given by [23], where the boundary conditions are enforced automatically by multiplying the neural network with a “distance function” $D(x)$. For our one-dimensional problem with homogeneous Dirichlet boundaries, the solution is thus represented as
@@
where $\sin(\pi x)$ is the distance function. This same discretization can be used in the “Deep Ritz” formulation of the problem [8], which obtains a numerical solution through minimizing the variational energy of Eq. (2) in terms of the neural network parameters.
Whether physics-informed or not, the choice of activation function plays a fundamental role in the training and performance of a neural network. Some common choices of activation functions are hyperbolic tangent ($\tanh(\bullet)$), sigmoid ($1/(1+\exp(-\bullet))$), softplus ($\ln(1+\exp(\bullet))$), and the rectified linear unit (ReLU) ($\max(0,\bullet)$). Other options include the swish activation and the Gaussian error linear unit (GELU) [12]. The problem of choosing activation functions has been studied extensively in the physics-informed machine learning literature. For example, in [27], using non-standard “hybrid” activation functions was shown to reduce errors in PINN solutions. Similarly, [15] showed that allowing the activation functions themselves to be partially learnable accelerated convergence in the physics-informed training. Another study found that learning the activation function allowed the network to better adapt to characteristics of the solution field [25]. For high-frequency solution fields, modifications inspired by the SIREN activation [21] were shown to reduce the number of parameters and thus expedite training [9]. In a comparative study on the Helmholtz equation, the swish activation function outperformed other activation functions by obtaining the smallest error with a reference solution [1]. In another work, the one-dimensional wave equation was used to compare the convergence and errors obtained from different activation functions, including hyperbolic tangent, swish, and sinusoidal activation functions, finding that no single activation performed best in all numerical experiments [17]. The authors remark that their choice to use the strong form of the governing equation places restrictions on the set of available activation functions. Namely, the governing equation requires second derivatives of the solution field, and the second derivative of ReLU is zero everywhere except at the origin. As such, they do not include ReLU in their study. Finally, [7] compared six different activation functions including hyperbolic tangent, ReLU, swish, and modifications thereof on incompressible flow problems, finding that ReLU consistently performed the worst.
@@
We use ADAM optimization with a learning rate of $1\times 10^{-3}$ to minimize the variational energy. Integration is done on a uniform grid of $250$ points and we run the optimization for $5000$ epochs. The exact solution is obtained by inspection as $\frac{100}{9\pi^{2}}\sin(3\pi x)$. As an additional point of reference, we train a second single hidden-layer network of the same width with hyperbolic tangent activation functions. See Figure 1 for the results. The converged energy of the ReLU network is notably larger than that of the network using hyperbolic tangent activation functions. The eyeball norm suffices to show that the ReLU network fails to reproduce the simple solution field. As the figure shows, this failure occurs despite the fact that the network can represent the exact solution when explicitly trained to do so. Figure 2 shows similar results for the higher-frequency source term $b(x)=100\sin(4\pi x)$. Evidently, these failures are not issues of expressiveness. We remark that the jagged shape of the solution field is similar to results obtained in [7], where ReLU provides very poor approximations of solution fields from incompressible flow problems. If the ReLU network is expressive enough to represent the true solution, and it possesses the same continuity properties as finite element discretizations, why does it perform so poorly with the Deep Ritz method? This is the question we address in the subsequent section.
@@
Figure 2: Training convergence of the variational energy objective with a higher-frequency source term for the two networks (left). The Deep Ritz solution with ReLU activation again provides an unphysical and inaccurate representation of the solution (center). We verify that the network is sufficiently expressive to represent the desired solution field (right).
@@ 2 Diagnosing the failure of ReLU in the Deep Ritz problem
where the discretized solution is given by Eq. (4). This gradient is used in the training process to iteratively update the parameters, but is often “out of sight” for the analyst, as it is computed by automatically differentiating the forward evaluation of the energy. We now compute each of the terms involved in Eq. (5) analytically. Recall that the solution field is given by the single hidden-layer network as
@@
Notice that second derivatives of ReLU appear in Eq. (8). This is a consequence of the nonlinearity of the discretization—while the $C^{0}$ continuity of the solution field is appropriate for a traditional Ritz method, training the network with the Deep Ritz method requires second derivatives of the activation. This is because trainable parameters lie within the ReLU activation, unlike a so called “linear” discretization, where the parameters only multiply spatial basis functions. Mathematically speaking, the delta function arising from the second derivative of ReLU is acceptable given that it appears under the integral in the variational energy. However, it is not immediately clear how PyTorch’s automatic differentiation handles gradients of discontinuous functions. We will investigate this numerically. To do this, we compute derivatives of the function $\sin(\pi x)\max(0,x-1/2)$ both analytically and with automatic differentiation. This function is reminiscent of one of the $N$ learned “basis” functions that are scaled and superimposed in the ReLU network of Eq. (4). To distinguish between the two cases of analytic and automatic differentiation, we call the analytically differentiated function $u(x)$ and the automatically differentiated function $w(x)$. Note that we can use Eqs. (6)-(8) with $N=1$, $\theta^{1}=1$, $\theta^{2}=-1/2$, and $\theta^{3}=1$ to obtain the relevant derivatives. To numerically approximate the delta function on a fixed integration grid, we require that
@@
where the notation $\mathbf{1}_{i}$ indicates a vector of zeros with a single $1$ in the $i$-th index. We take the finite difference step size to be $h=1\times 10^{-3}$, the learning rate to be $5\times 10^{-3}$, and run ADAM optimization for $5000$ epochs. The first single hidden-layer ReLU network is trained with gradients from automatic differentiation (AD) and the second is trained with manually supplied gradients computed using finite differences (FD). See Figure 6 for the results of this comparison. The network with finite differenced gradients obtains a lower converged energy value, and a smoother, more accurate solution field.
@@ 3 Conclusion
The purpose of this short work has been to explain failures of ReLU activation functions for physics-informed machine learning problems. Though the objective in the Deep Ritz test problem we considered required only first derivatives of the solution field, we showed that training the network requires second derivatives of the activation function. The second derivatives of the $C^{0}$ ReLU activation give rise to delta functions which automatic differentiation does not account for, thus leading to inaccurate gradient computations. We obtained accurate gradient computations by finite differencing the loss function, and showed that this remedied the failures of the ReLU network on the test problem. Though the corrected gradients improved performance, we then showed that the loss landscape of our test problem with the ReLU network is extremely noisy, and that the noise promptly vanishes when higher-order continuity is introduced. This study suggests that ReLU activation is a poor choice for physics-informed machine learning, which corroborates the conclusions of past comparisons of activation functions in the PINNs literature.